A businessman borrows P20,000 for his operations. One financial institution is willing to lend him money
for one year at a discounted interest rate of 10.5% per annum. What is the effective rate?
1
Expert's answer
2021-09-23T03:03:55-0400

Here, P = 1000, r = 0.08, and m = 1. 

  Thus, i = r = 0.08 and n = 3, so

 "A=P(1+r\/m)^n"

"=1000(1+0.08\/1)^3"

"=1000(1.08)^3"

=1259.71

or $1259.71.

once again but m value is changed

Here, P = 1000, r = 0.08, and m = 2. 

  Thus,  and n = (3)(2) = 6, so


 "A=P(1+r\/m)^n"

"=1000(1+0.08\/2)^3"

"=1000(1.04)^3"

  $1265.32.